# Frobenius intertwiners for $q$-difference equations

### v1

@@ 2. Vertex of $X=T^{*}\mathbb{P}^{n-1}$
where $S^{a_{i}}(\boldsymbol{a},h,q,z)$ and $S^{h}(\boldsymbol{a},h,q,z)$ denote matrices whose entries are rational functions in the equivariant parameters ${\boldsymbol{u}}$, $\hbar$, $q$ and $z$:
$S^{a_{i}}(\boldsymbol{a},h,q,z),S^{h}(\boldsymbol{a},h,q,z)\in GL_{n}(\mathbb{Q}({\boldsymbol{u}},\hbar,z,q)).$
3. $p$adic $q$-hypergeometric functions
3.1.
Let $p$ be a prime, $\mathbb{Q}_{p}$ be the filed of $p$-adic numbers, $\mathbb{Z}_{p}$ be the integers of $\mathbb{Q}_{p}$. Starting from this section $\mathbb{K}$ denotes the $p$-adic completion of the algebraic closure of $\mathbb{Q}_{p}$. Let us fix $q\in\mathbb{K}^{\times}$ of the following form
(3.1) $q=1+t,\ \ t\neq 0,\ \ |t|_{p}\leqslant 1.$
From now, we assume that all the equivariant parameters (1.14) are $p$ adic integers $a_{1},\dots,a_{n},h\in\mathbb{Z}_{p}$.
3.2.
@@ 3. $p$adic $q$-hypergeometric functions
Let $\overline{\mathbb{K}(z)}$ be the completion of $\mathbb{K}(z)$ with respect to the Gauss norm:
$\left|\sum\limits_{m=0}^{\infty}\,r_{m}(z)\right|_{\mathrm{Gauss}}=\ \ \max\,\{\,|r_{m}(z)|_{p}:m=0,1,2,\dots\}.$
@@
A Frobenius intertwiner from $\mathcal{H}(\boldsymbol{a},h,q^{p},z^{p})$ to $\mathcal{H}(\boldsymbol{a}^{\prime},h^{\prime},q,z)$ is $\mathsf{U}(z)\in\mathsf{GL}_{n}(\overline{\mathbb{K}(z)})$ such that
$\mathbf{M}_{\mathcal{L}}(\boldsymbol{a}^{\prime},\hbar^{\prime},q,z)=\mathsf{U}(zq)^{-1}\,\mathbf{M}_{\mathcal{L}}(\boldsymbol{a},\hbar,q^{p},z^{p})\,\mathsf{U}(z).$
@@
If ${\Psi}(\boldsymbol{a},\hbar,q^{p},z^{p})$ is a fundamental solution matrix for the hypergeometric system $\mathcal{H}(\boldsymbol{a},\hbar,q^{p},z^{p})$ (1.11) then $\mathsf{U}(z){\Psi}(\boldsymbol{a},\hbar,q^{p},z^{p})$ is a fundamental solution matrix for $\mathcal{H}(\boldsymbol{a}^{\prime},h^{\prime},q,z)$. Thus, it is of the form ${\Psi}(\boldsymbol{a}^{\prime},\hbar^{\prime},q,z)\Lambda$ and we obtain a relation
(3.2) ${\Psi}(\boldsymbol{a}^{\prime},\hbar^{\prime},q,z)\Lambda=\mathsf{U}(z){\Psi}(\boldsymbol{a},\hbar,q^{p},z^{p})$
@@
for some constant matrix $\Lambda\in\mathsf{GL}_{n}(\mathsf{\mathbb{K}})$. The matrix $\mathsf{U}(0)$ is called the constant term of the Frobenius intertwiner. If a fundamental solution is normalized by ${\Psi}(\boldsymbol{a},\hbar,q,0)=1$, then $\mathsf{U}(0)=\Lambda$.
3.3.
For the purposes of this note, it is sufficient to restrict to a smaller ring $\widehat{\mathsf{R}}_{p}\subset\overline{\mathbb{K}(z)}$ which has a more explicit structure. Let $\mathsf{R}_{p}=\mathbb{Z}[z,(1-z)^{-1}]$ be the ring of rational functions with denominator given by powers of $(1-z)^{-1}$. The ring we need is the $p$-adic completion
$\widehat{\mathsf{R}}_{p}=\lim\limits_{\leftarrow}\,\mathsf{R}_{p}/p^{s}\mathsf{R}_{p}.$
By definition, the elements of $\widehat{\mathsf{R}}_{p}$ are sums of the form
$\sum\limits_{k=0}^{\infty}\,p^{k}\dfrac{g_{k}(z)}{(1-z)^{m_{k}}}$
where $m_{k}\in\mathbb{N}$ and $g_{k}(z)\in\mathbb{Z}[z]$. Put differently, an element $g(z)\in\widehat{\mathsf{R}}_{p}$ can be represented by a power series $g(z)\in\mathbb{Z}[[z]]$ such that for any $s\geqslant 0$ there is a natural number $m_{s}$ such that
$(1-z)^{m_{s}}g(z)\!\!\pmod{p^{s+1}}\in\mathbb{Z}[z].$
This condition provides a useful numerical test to determine whether a given power series $g(z)$ belongs to $\widehat{\mathsf{R}}_{p}$ or not, see Section 5. The Frobenius intertwiners we encounter in this paper are all $\mathsf{GL}_{n}(\widehat{\mathsf{R}}_{p})$ valued.
3.4.
For $n\in\mathbb{N}$ we define the $q$-number by:
$[n]_{q}=\dfrac{1-q^{n}}{1-q}=1+q+\dots+q^{n-1}.$
Let $[n]_{q}!=[1]_{q}[2]_{q}\dots[n]_{q}$ denote the $q$-factorial and
(3.3) $\binom{n}{k}_{\!\!q}=\dfrac{[n]_{q}!}{[k]_{q}![n-k]_{q}!}$
be the $q$-binomial factor. We recall the original definitions of the $p$-adic gamma and $q$-gamma functions:
Theorem 3.1 (Koblitz, [Ko80]).
For $q$ as in (3.1) the function $\Gamma_{p,q}(n)=(-1)^{n}\prod\limits_{{0<i<n}\atop{p\nmid i}}\,[n]_{q}$ extends to a continuous function on $\mathbb{Z}_{p}$. $\lim\limits_{q\to 1}\,\Gamma_{p,q}=\Gamma_{p}$ where $\Gamma_{p}$ denotes the $p$-adic gamma function of Morita [Mo75], which is the extension to $\mathbb{Z}_{p}$ of $\Gamma_{p}(n)=(-1)^{n}\prod\limits_{{0<i<n}\atop{p\nmid i}}\,n.$
3.5.
We formulate our main theorem in the geometric language of Section 2. For this, we replace the equivariant K-theory (2.1) by
(3.4) $K_{T}(X,\mathbb{K})=K_{T}(X)\otimes\mathbb{K}=\mathbb{K}[L^{\pm},u_{1}^{\pm},\dots,u^{\pm}_{n},\hbar^{\pm}]/(L-u_{1})\dots(L-u_{n}).$
The Frobenius intertwiner is understood as a map
$\mathsf{U}(z):K_{T}(X,\mathbb{K})[[z]]\longrightarrow K_{T}(X,\mathbb{K})[[z]],$
In particular, the constant term acts on $K$-theory
$\mathsf{U}(0):K_{T}(X,\mathbb{K})\longrightarrow K_{T}(X,\mathbb{K}).$
We are now ready to formulate our main result:
@@
For any $\boldsymbol{a}=(a_{1},\dots,a_{n})\in\mathbb{Z}^{n}_{p},\ \ h\in\mathbb{Z}_{p}$ there exists unique Frobenius intertwiner $\mathsf{U}(z)\in\mathsf{GL}_{n}(\overline{\mathbb{K}(z)})$ from $\mathcal{H}(\boldsymbol{a},h,q^{p},z^{p})$ to $\mathcal{H}(p\boldsymbol{a},ph,q,z)$, where $p\boldsymbol{a}=(pa_{1},\dots,pa_{n})$. The constant term of the intertwiner acts in $K_{T}(X,\mathbb{K})$ as the operator of multiplication by the K-theory class
(3.5) $\mathsf{U}(0)=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p,q}(px+ph-pa_{j})}{\Gamma_{p,q}(px-pa_{j})\Gamma_{p,q}(ph)}.$
Clearly, multiplying a Frobenius intertwiner by a constant multiple $\mathsf{U}(z)\to c\mathsf{U}(z)$, $c\in\mathbb{K}^{\times}$ preserves all the properties. So, the “uniqueness” in Theorem 3.2 should be understood as uniqueness up to a scalar multiple.
If the torus fixed points of $T$-action are isolated, multiplication by any equivariant K-theory class is diagonal in the basis given by the classes of $T$-fixed points. In our case $X=T^{*}\mathbb{P}^{n-1}$ the $X^{T}$ is a finite set when all $a_{1},\dots,a_{n}$ are pairwise distinct. The corresponding fixed point component of $\mathsf{U}(0)$ are obtained by specialization $x=a_{i}$. We thus conclude with the following result.

### v2

@@ 3. Vertex function of $X=T^{*}\mathbb{P}^{n-1}$
where $A_{i}(\boldsymbol{a},h,q,z)$ and $H(\boldsymbol{a},h,q,z)$ denote matrices whose entries are rational functions in the equivariant parameters ${\boldsymbol{u}}$, $\hbar$, $q$ and $z$: $A_{i}(\boldsymbol{a},h,q,z),H(\boldsymbol{a},h,q,z)\in\mathsf{GL}_{n}(\mathbb{Q}({\boldsymbol{u}},\hbar,z,q)).$
These difference equations are useful for the analysis of QDEs, particularly when the parameters $h$, $\boldsymbol{a}$ are specialized to $\mathbb{Z}_{p}$-values, to which we now switch out attention.
4. $p$ adic completions and norms
In this section, we provide a brief overview of $p$-adic analysis, $p$-adic analytic functions, and give several illustrative examples of simple Frobenius structures. Our running example in this section is the quantum differential and quantum difference equation associated with a “point” $X=T^{*}\mathbb{P}^{0}$. The Frobenius structures arising in this case can be described using only elementary combinatorics.
4.1.
We denote by $|\cdot|_{p}$ the $p$-adic norm on $\mathbb{Q}$ normalized so that $|p^{s}|=1/p^{s}$. The field of $p$-adic numbers $\mathbb{Q}_{p}$ is the completion of $\mathbb{Q}$ with respect to this norm. $\mathbb{Z}_{p}\subset\mathbb{Q}_{p}$ denotes the set of integers of $\mathbb{Q}_{p}$ which is the subset of elements with norm $|x|_{p}\leqslant 1$. We also denote by $\Omega$ the completion of the algebraic closure of $\mathbb{Q}_{p}$.
Let $f(z)$ be a rational function:
$f(z)=\dfrac{\sum_{i}a_{i}z^{i}}{\sum_{j}b_{j}z^{j}}\in\mathbb{Q}(z)$
The $p$-adic Gauss norm on $\mathbb{Q}(z)$ is defined by
$|f(z)|_{\rm Gauss}:=\dfrac{\textrm{max}_{i}\{|a_{i}|_{p}\}}{\textrm{max}_{j}\{|b_{i}|_{p}\}}$
@@ 4. $p$ adic completions and norms
We denote by $E_{p}$ the field of $p$-adic analytic functions which is the completion of $\mathbb{Q}(z)$ with respect to the Gauss norm. By definition, an element $f(z)\in E_{p}$ is a sum
$f(z)=\sum\limits_{n=0}^{\infty}\,f_{i}(z),\ \ f_{i}(z)\in\mathbb{Q}(z)$
converging in the Gauss norm. Note that $|f(z)|_{\rm Gauss}=\max\{|f_{i}(z)|_{\rm Gauss}:i=0,1,\dots\}.$
4.2.
One of the main questions we encounter in this paper is the following: let $f(z)\in\mathbb{Q}[[z]]$ be a power series. When is $f(z)$ a Taylor expansion of an element $g(z)\in E_{p}$?
Since the norm of analytic elements is bounded, there exist $n$ such that $|p^{n}g(z)|_{\rm Gauss}\leqslant 1$. Thus, if $f(z)$ is the series expansion of $g(z)\in E_{p}$ then we can “clear denominators” - the coefficients of power series $p^{n}f(z)$ do not have powers of $p$ appearing in the denominators. In particular, the coefficients of the power series $p^{n}f(z)$ have well defined reductions $\pmod{p^{s}}$ for $s=1,2,\dots$. These reductions can be used to determine whether $f(z)$ is the expansion of an analytic function using the following elementary lemma.
Lemma 4.1.
A power series $f(z)\in\mathbb{Q}[[z]]$ is a Taylor expansion of an analytic element $g(z)\in E_{p}$ if and only if there exist $n\in\mathbb{N}$ such that for any $s\in\mathbb{N}$
(4.1) $p^{n}f(z)\equiv g_{s}(z)\pmod{p^{s}}$
where $g_{s}(z)$ is a power series expansion of some rational function from $\mathbb{Q}(z)$.
Proof.
First assume $f(z)\in\mathbb{Q}[[z]]$ is such that (4.1) holds for some $n$. Let us define $f_{i}(z)$ by
$f_{i}(z)=g_{i+1}(z)-g_{i}(z)$
By construction $|f_{i}(z)|_{p}<1/p^{i}$, thus the sum $\sum_{i=0}^{\infty}f_{i}(z)$ converges in the Gauss norm and we obtain:
$f(z)=\dfrac{1}{p^{n}}\sum_{i=0}^{\infty}f_{i}(z)\in E_{p}$
For the opposite direction, assume $f(z)$ is a power series in $z$ which is a Taylor expansion of some ${g}(z)\in E_{p}$. We have
(4.2) ${g}(z)=\sum\limits_{i=0}^{\infty}\,g_{i}(z)$
for some $g_{i}(z)\in\mathbb{Q}(z)$. Assume that $|{g}(z)|_{\rm Gauss}=p^{n}$ for some $n\in\mathbb{Z}$. Since (4.2) converges in the Gauss norm it follows that for any $s\in\mathbb{N}$ there is $N_{s}\in\mathbb{N}$ such that $p^{n}g_{i}(z)\equiv 0\pmod{p^{s}}$ for all $i>N_{s}$. Therefore,
$p^{n}f(z)\equiv\sum\limits_{n=0}^{N_{s}}\,g_{i}(z)\pmod{p^{s}}$
In Section 8 we will demonstrate how this Lemma can be used to discover non-trivial Frobenius structures numerically.
Example: Let $h\in\mathbb{Q}$ be a rational number such that $h\in\mathbb{Z}_{p}$. Let us consider a power series
(4.3) $f(h,z)=\sum\limits_{n=0}^{\infty}\,(-1)^{n}\binom{h}{n}z^{n}$
which is a Taylor expansion of $(1-z)^{h}$. It is known that this series is not an expansion of an element from $E_{p}$ for generic $h$. However, the power series
(4.4) $\mathsf{U}(h,z)=\dfrac{f(ph,z)}{f(h,z^{p})}$
is in $E_{p}$. Indeed
$\mathsf{U}(h,z)=\dfrac{(1-z)^{ph}}{(1-z^{p})^{h}}=\Big{(}1+\dfrac{(1-z)^{p}-(1-z^{p})}{1-z^{p}}\Big{)}^{h}$
Note that $(1-z)^{p}-(1-z^{p})=p\alpha(z)$ for some $\alpha(z)\in\mathbb{Z}[z]$ and therefore
$\mathsf{U}(h,z)=\sum\limits_{n=0}^{\infty}\,\binom{h}{n}p^{n}\dfrac{\alpha(z)^{n}}{(1-z^{p})^{n}}$
For $h\in\mathbb{Z}_{p}$ we also have $\binom{h}{n}\in\mathbb{Z}_{p}$ for any $n$. From which we see that:
(4.5) $\mathsf{U}(h,z)\equiv\sum\limits_{n=0}^{s-1}\,\binom{h}{n}p^{n}\dfrac{\alpha(z)^{n}}{(1-z^{p})^{n}}\pmod{p^{s}}$
Therefore $\mathsf{U}(h,z)\in E_{p}$ by the previous Lemma.
4.3.
Let $V(\alpha,z)$ denote the finite dimensional $\mathbb{Q}_{p}$-space of solutions to some $p$-adic differential equation in $z$ with parameters $\alpha$. As we explain in the next section, a Frobenius intertwiner gives a map:
$V(\alpha,z)\stackrel{{\scriptstyle\mathsf{U}(z)}}{{\longrightarrow}}V(\alpha^{\prime},z^{p})$
As a map between two different vector spaces $\mathsf{U}(z)$ carries no interesting information. However, when $\alpha^{\prime}=\alpha$ at the points $z^{p}=z$ it becomes an automorphism of $V(\alpha,z)$ and the characteristic polynomial of $\mathsf{U}(z)$ at such points is an interesting invariants of the QDEs. The invariants appearing in this way find many applications in the theory of Gauss and exponential sums [Dw74, Sp80, K90].
The points $z\in\mathbb{Q}_{p}$ satisfying $z^{p}=z$ are the Teichmüller elements: recall that for any $a\in\mathbb{F}_{p}$ there is unique $[a]\in\mathbb{Q}_{p}$ such that $[a]^{p}=[a]$ and $[a]\equiv a\pmod{p}$. Clearly, the map $\mathbb{F}_{p}^{\times}\to\mathbb{Q}^{\times}_{p}$, sending $a$ to $[a]$, is a multiplicative character. The element $[a]\in\mathbb{Q}_{p}$ is called the Teichmüller representative or the Teichmüller lift of $a$.
By definition, if $a\neq 0$ then $|[a]|_{p}=1$. For this reason, a power series $f(z)\in\mathbb{Q}[[z]]$ usually can not be evaluated at such points, since the convergence is not guaranteed. A notable property of the analytic elements $g(z)\in E_{p}$ is that they often can be evaluated at points $z\in\mathbb{Z}_{p}$ with $|z|_{p}=1$: if $f(z)\in E_{p}$ then up to terms with norm $1/p^{s}$ it can be approximated by a rational function. The values of these rational functions at $[a]$ then approximate $f([a])$ with an error of norm $1/p^{s}$.
For these reasons, Frobenius intertwiners are naturally required to have coefficients in $E_{p}$. This motivates the definition we give in the next section.
Example: Continuing our running example, we note that the power series (4.3) spans $1$-dimensional space of solutions of the simplest hypergeometric differential equation
(4.6) $(1-z)z\dfrac{df(z)}{dz}+zhf(z)=0$
In the hypergeometric notations $f(z)={}_{0}F_{1}(-h;;z)$. By its definition, multiplication by power series (4.4) maps solutions with parameters $(h,z^{p})$ to solutions with parameters $(ph,z)$. As we have already seen, $U(h,z)$ is in $E_{p}$. Thus, $U(h,z)$ is a Frobenius intertwiner between these solutions.
Assume that $h$ is such that $(p-1)h\in\mathbb{Z}$. Then $(1-z)^{h(1-p)}$ is a rational function and therefore the composition
(4.7) $\mathsf{U}^{\prime}(z)=(1-z)^{h(1-p)}\mathsf{U}(h,z)=\dfrac{(1-z)^{h}}{(1-z^{p})^{h}}$
is also in $E_{p}$. Thus $\mathsf{U}^{\prime}(z)$ is a Frobenius intertwiner between solutions at $(h,z^{p})$ and $(h,z)$. Using the approximations by rational functions (4.5) one can prove that the value of this intertwiner at the Teichmüller elements equals:
(4.8) $\mathsf{U}^{\prime}([a])=[a-1]^{h(1-p)}$
Note that a naive substitution of $z=[a]$ to (4.7) would give an incorrect answer $\mathsf{U}^{\prime}([a])=1$ (since $[a]^{p}=[a]$). The $p$-adic unit root (4.8) is the “interesting invariant” of the differential equation (4.6) which we mentioned above.
4.4.
Let show that $\mathsf{U}(h,z)$ given by (4.4) is in $E_{p}$ in a different way using the $q$-deformation. First, let us fix $q\in\Omega$ of the form $q=1+t$ with $|t|_{p}<1$. In this case, for $h\in\mathbb{Z}_{p}$ we have well defined elements $q^{\pm h}\in\Omega$. The $q$-deformation of the hypergeometric $f(z)={}_{0}F_{1}(-h;;z)$ is given by the power series
(4.9) $f(h,q,z)=\sum\limits_{n=0}^{\infty}\,\dfrac{(q^{-h},q)_{n}}{(q,q)_{n}}z^{n},\ \ \ (x,q)_{n}=(1-x)(1-xq)\dots(1-xq^{n-1}),$
which solves the $q$-analog of (4.6):
(4.10) $f(h,q,zq)=\dfrac{(1-z)}{(1-zq^{-h})}f(h,q,z)$
At $h\in\mathbb{N}$ this power series truncates to a Laurent polynomial
(4.11) $f(h,q,z)=(1-zq^{-1})(1-zq^{-2})\dots(1-zq^{-h}).$
Assume that $\zeta\neq 1$ is a unit root $\zeta^{p^{s}}=1$. Assume further that the order of the root $\zeta$ is $p^{m}$ for some $m\leqslant s$. Let $h\in\mathbb{N}$ and write $h=h^{(s)}+p^{s-1}h^{\prime}$ for some natural numbers $h^{(s)}$ and $h^{\prime}$. From (4.11) we find
$f(ph,\zeta,z)=f(ph^{(s)},\zeta,z)(1-z^{p^{m}})^{h^{\prime}p^{s-m}}$
Since $\zeta^{p}$ is also a unit root or order $p^{m-1}$ we obtain similarly
$f(h,\zeta^{p},z^{p})=f(h^{(s)},\zeta^{p},z^{p})(1-z^{p^{m}})^{h^{\prime}p^{s-m}}$
where in the last two equalities we used that $\prod_{i=0}^{m-1}(1-z\zeta^{i})=(1-z^{m})$ if $\zeta$ has order $m$.
Let us consider the power series
$\mathsf{U}(h,q,z)=\dfrac{f(ph,q,z)}{f(h,q^{p},z^{p})}$
From the above computations, we see that for $h\in\mathbb{N}$ we have
(4.12) $\mathsf{U}(h,\zeta,z)=\mathsf{U}(h^{(s)},\zeta,z)$
where $h^{(s)}$ is a natural number defined by $h^{(s)}=h\pmod{p^{s-1}}$.
The equation (4.12) is crucial. Note that (4.12) extends to $p$-adic integers $h\in\mathbb{Z}_{p}$. Clearly, unit roots $q^{p^{s}}=1$ such that $q\neq 1$ are the roots of the polynomial:
$[p^{s}]_{q}=\dfrac{1-q^{p^{s}}}{1-q}$
Thus, the relation (4.12) implies that for any $h\in\mathbb{Z}_{p}$ we have
$\mathsf{U}(h,q,z)=\mathsf{U}(h^{(s)},q,z)+[p^{s}]_{p}\,\Big{(}\ \ \dots\ \ \Big{)}$
where $\mathsf{U}(h^{(s)},q,z)$ is a rational function (since $h^{(s)}$ is integral) and $\dots$ stands for a power series in $z$ with coefficients regular at $p^{s}$-th roots. Note that for $q=1+t$ with $|t|_{p}<1$ we have $[p^{l}]_{p}=0\pmod{p^{s}}$ for sufficiently large $l$ thus
(4.13) $\mathsf{U}(h,q,z)\pmod{p^{s}}.$
is a rational function of $z$.
In the limit $q\to 1$ $\mathsf{U}(h,q,z)$ converges to (4.4) and therefore $\mathsf{U}(h,z)\pmod{p^{s}}.$ is also rational and by Lemma 4.1 we conclude that $\mathsf{U}(h,z)\in E_{p}$.
The above computation (4.13) also shows what $\mathsf{U}(h,q,z)\in\hat{E}_{p}$ where $\hat{E}_{p}$ is the completion of $\Omega(z)$ with respect to Gauss norm (we need to replace $\mathbb{Q}$ with $\Omega$ to accommodate for $q\in\Omega$). Thus, $\mathsf{U}(h,q,z)$ is a Frobenius intertwiner between $f(h,q^{p},z^{p})$ and $f(ph,q,z)$.
Finally, let us note that the hypergeometric differential equation (4.6) is the quantum differential equation associated with a zero-dimensional quiver variety $X=T^{*}\mathbb{P}^{0}$ (enumerative geometry of such varieties may be non-trivial, see [DS]). The $q$-difference equation (4.10) is the $K$-theoretic quantum difference equation for this $X$. The power series (4.3) and (4.9) represent the “fundamental solution matrices” of these equations. The running example of this section is thus, the most basic $n=1$ case of the hypergeometric functions (2.3) we consider in this paper.
5. Frobenius intertwiner for $q$-hypergeometric equation
5.1.
Let us consider the $q$-difference hypergeometric equation (2.10). Unless otherwise specified, we assume throughout this section that $h\in\mathbb{Z}_{p}$, $\boldsymbol{a}\in\mathbb{Z}^{n}_{p}$. We assume that $q\in\Omega$ is of the form $q=1+t$ with $|t|_{p}<1$. In this case
$\hbar=q^{h},\ \ u_{1}=q^{a_{1}},\dots,u_{n}=q^{a_{n}}$
are well defined in $\Omega$. Let $\hat{E}_{p}$ be the completion of $\Omega(z)$ with respect to the Gauss norm. The following is motivated by discussion of Section 4.
Definition 5.1.
@@ 5. Frobenius intertwiner for $q$-hypergeometric equation
A Frobenius intertwiner from a $q$-hypergeometric system with parameters $(h,\boldsymbol{a},q^{p},z^{p})$ to a $q$-hypergeometric system with parameters $(h^{\prime},\boldsymbol{a}^{\prime},q,z)$ is an element $\mathsf{U}(z)\in\mathsf{GL}_{n}(\hat{E}_{p})$ such that $\mathbf{M}(\hbar,\boldsymbol{a},q^{p},z^{p})=\mathsf{U}(zq)^{-1}\,\mathbf{M}(\hbar^{\prime},\boldsymbol{a}^{\prime},q,z)\,\mathsf{U}(z)$.
@@
If ${\Psi}(h,\boldsymbol{a},q^{p},z^{p})$ is a fundamental solution matrix for the $q$-hypergeometric system with parameters $(h,\boldsymbol{a},q^{p},z^{p})$ and $\mathsf{U}(h,\boldsymbol{a},q,z)$ is a Frobenius intertwiner as in the Definition above then $\mathsf{U}(h,\boldsymbol{a},q,z){\Psi}(h,\boldsymbol{a},q^{p},z^{p})$ is a fundamental solution matrix for the $q$-hypergeometric system with parameters $(h^{\prime},\boldsymbol{a}^{\prime},q,z)$. Thus, it is of the form ${\Psi}(h^{\prime},\boldsymbol{a}^{\prime},q,z)\Lambda$ where $\Lambda\in\mathsf{GL}_{n}(\Omega)$ is some constant matrix. We obtain a relation
(5.1) ${\Psi}(h^{\prime},\boldsymbol{a}^{\prime},q,z)\Lambda=\mathsf{U}(h,\boldsymbol{a},q,z){\Psi}(h,\boldsymbol{a},q^{p},z^{p}).$
@@
$\mathsf{U}(h,\boldsymbol{a},q,0)$ is called the constant term of the Frobenius intertwiner. If the fundamental solution matrix is normalized by ${\Psi}(h,\boldsymbol{a},q,0)={\rm Id}$ (the identity matrix), then $\mathsf{U}(0)=\Lambda$.
5.2.
Let us consider the power series:
(5.2) $\mathsf{U}(h,\boldsymbol{a},q,z)=\tilde{\Psi}(ph,p\boldsymbol{a},q,z)\tilde{\Psi}(h,\boldsymbol{a},q^{p},z^{p})^{-1}\in K_{T}(X)^{\otimes 2}[[z]]$
where $\tilde{\Psi}(h,\boldsymbol{a},q,z)$ denotes the fundamental solution matrix normalized as in (3.5). Using the natural non-degenerate paring on $K_{T}(X)$ we can also view it as an operator
$\mathsf{U}(h,\boldsymbol{a},q,z):K_{T}(X)[[z]]\longrightarrow K_{T}(X)[[z]]$
In particular, the constant term is a linear map in $K$-theory
(5.3) $\mathsf{U}(h,\boldsymbol{a},q,0):K_{T}(X)\longrightarrow K_{T}(X)$
The goal of this section is to prove the following result (see Section 3.1 for notations and description of the ring $K_{T}(X)$):
@@
The power series $\mathsf{U}(h,\boldsymbol{a},q,z)$ defined by (5.2) is a Frobenius intertwiner between the $q$-hypergeometric systems with parameters $(h,\boldsymbol{a},q^{p},z^{p})$ and $(ph,p\boldsymbol{a},q,z)$, where $p\boldsymbol{a}=(pa_{1},\dots,pa_{n})$. The constant term of the intertwiner (5.3) acts in $K_{T}(X)$ as the operator of multiplication by the K-theory class
(5.4) $\mathsf{U}(h,\boldsymbol{a},q,0)=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p,q}(px+ph-pa_{j})}{\Gamma_{p,q}(px-pa_{j})\Gamma_{p,q}(ph)}.$
Definition of $p$-adic $q$-Gamma function $\Gamma_{p,q}$ is reminded in Section A.3.
5.3.
Let us first analyze the constant term of this operator.
Proposition 5.3.
The constant term of the power series (5.2) acts in $K_{T}(X)$ as multiplication by the following K-theory class:
(5.5) $\mathsf{U}(h,\boldsymbol{a},q,0)=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p,q}(px+ph-pa_{j})}{\Gamma_{p,q}(px-pa_{j})\Gamma_{p,q}(ph)}$
where $\Gamma_{p,q}$ denotes the $p$-adic $q$-gamma function .
Proof.
From the normalization (3.5) we find:
$\mathsf{U}(h,\boldsymbol{a},q,0)=\dfrac{\Phi(ph,p\boldsymbol{a},px,q,z)}{\Phi(h,\boldsymbol{a},x,q^{p},z^{p})}$
From definition (3.2) we see that this ratio is exactly the product of factors as in (A.7). The Proposition follows by Lemma A.4. ∎
If the set of the $T$-fixed points $X^{T}$ is finite, then multiplication by any equivariant $K$-theory class is diagonal in the basis of $K_{T}(X)$ given by the classes of $T$-fixed points. In our case $X=T^{*}\mathbb{P}^{n-1}$ the set $X^{T}$ is finite when all $a_{1},\dots,a_{n}$ are pairwise distinct. The corresponding eigenvalues of multiplication by $\mathsf{U}(h,\boldsymbol{a},q,0)$ are obtained by specialization $x=a_{i}$. We thus conclude with the following result.